# Feature Importance-Aware Deep Joint Source-Channel Coding for Computationally Efficient and Adjustable Image Transmission

### v1

@@ Abstract
Recent advancements in deep learning-based joint source-channel coding (deepJSCC) have significantly improved communication performance, but their high computational demands restrict practical deployment. Furthermore, some applications require the adaptive adjustment of computational complexity. To address these challenges, we propose a computationally efficient and adjustable deepJSCC model for image transmission, which we call feature importance-aware deepJSCC (FAJSCC). Unlike existing deepJSCC models that equally process all neural features of images, FAJSCC first classifies features into important and less important features and then processes them differently. Specifically, computationally-intensive self-attention is applied to the important features and computationally-efficient spatial attention to the less important ones. The feature classification is based on the available computational budget and importance scores predicted by an importance predictor, which estimates each feature’s contribution to performance. It also allows independent adjustment of encoder and decoder complexity within a single trained model. With these properties, our FAJSCC is the first deepJSCC that is computationally efficient and adjustable while maintaining high performance. Experiments demonstrate that our FAJSCC achieves higher image transmission performance across various channel conditions while using less computational complexity than the recent state-of-the-art models. Adding to this, by separately varying the computational resources of the encoder and decoder, it is concluded that the decoder’s error correction function requires the largest computational complexity in FAJSCC, which is the first observation in deepJSCC literature. The FAJSCC code is publicly available at https://github.com/hansung-choi/FAJSCC.
@@ I Introduction
Despite its success in communication performance, existing deepJSCC research suffers from huge computational complexity incurred by large neural networks, along with associated problems like high power consumption, hardware costs, and latency, which prohibits its practical deployment. To reduce such computational burden, model compression techniques in deep learning are widely used. For instance, pruning [28, 29, 30] and low-rank decomposition [31] compress the trained models, by which they successfully reduce the number of model parameters (i.e., less computations) at the expense of performance degradation and extra training efforts. Another line of computationally-efficient works relies on better feature selection. Note that some features in data contain larger information (e.g., features about main objectives) while some contain less (e.g., about backgrounds) [32]. In the context of deepJSCC, [33] incorporates part of this idea by simply omitting the least important features in the computation. However, this approach also compromises task performance in order to reduce computational resources.
@@
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in surveillance applications, it is essential to transmit images of large surveillance areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for monitoring while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application level, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy. However, to the best of our knowledge, this issue is largely unstudied in existing literature.
To address both efficiency and dynamic control over computational complexity, we propose feature importance-aware deepJSCC, or FAJSCC for short. Similar to the approach of adopting different computation blocks for different feature importance in super-resolution deep learning research [34, 35], FAJSCC processes important features with detailed self-attention [36, 37, 38] and less important features with fast spatial attention [39]. To this end, we introduced a new feature importance-aware block (FA block) that differently processes features. In each FA block, there is a tiny importance predictor that scores patch-wise importance of feature. The importance score captures the efficiency of performance improvement when allocating higher computational resources to a given feature. In other words, features that yield greater performance improvements when assigned more computational resources receive higher importance scores. Using patch-wise importance scores, the FA block first classifies features into important and less important features. It then applies computationally intensive self-attention to important feature patches and computationally efficient spatial attention to less important ones.
To adapt to varying computational budgets, we introduce importance ratio $\gamma$, which indicates the fraction of important features. When the importance ratio is $1$, all patches are processed using the computationally intensive self-attention mechanism. When it is 00, none are processed by self-attention. By adjusting the importance ratio, we can control computational complexity, making FAJSCC the first deepJSCC model with adjustable computational complexity within a single model. Additionally, since FAJSCC can decouple computational complexity from the model architecture, it allows for observing the impact of the encoder’s and decoder’s complexity on performance. This is not possible with existing JSCC models where computational complexity is fixed at the encoder and decoder and fully tied to a trained model. Note that the encoder compresses the images and adds redundancy to protect the data from channel noise, while the decoder corrects errors and (re)generates the transmitted images. Then, one might wonder which of these functions demands the highest computational complexity, and under what conditions. By independently varying the importance ratios at the encoder and decoder across different signal-to-noise ratios (SNRs), we can make two key conclusions: 1) the decoder requires higher computational complexity than the encoder, and 2) the decoder’s error correction functionality demands even higher computational complexity when the SNR is low. This is the first result to provide a detailed analysis of computational complexity in deepJSCC.
As a consequence of our design, our FAJSCC achieves better image reconstruction accuracy than other deepJSCC models, even with smaller computational complexity compared to the previous state-of-the-art (SOTA) model, SwinJSCC [8]. This improvement is primarily due to the computational resource allocation enabled by importance-wise feature computation. Since FAJSCC processes less important features using computationally light spatial attention, it saves significant computational resources compared to SwinJSCC, which applies computationally heavy self-attention to all features. By reallocating these saved resources to enhance important feature processing, FAJSCC outperforms SwinJSCC even at lower computational complexity. Additionally, we designed FAJSCC to be robust against various communication channel SNRs by leveraging a two-step SNR-adaptive feature extraction.
Our main contributions can be summarized as follows.
  • Efficient Computation: Our FAJSCC efficiently allocates computational resources by distinguishing between important and less important features by importance scores. Important features are processed with computationally intensive self-attention, while less important ones are handled with computationally efficient spatial attention, reducing overall computational complexity. Compared to recent SwinJSCC, FAJSCC achieves better image reconstruction accuracy with lower computational complexity by reallocating saved resources to enhance important feature processing. Moreover, to train an SNR-adaptive model, FAJSCC incorporates a two-step SNR feature extraction to refine random SNR information for each processing block with minimal computational cost, making it robust to various communication channel noises.
  • Adjustable Computation: Our FAJSCC introduces a dynamic computational complexity control mechanism through an importance ratio, adjusting the number of important features. This adjustment can be applied separately to both the encoder and decoder without needing to agree with each other’s value. Also, FAJSCC maintains high performance across different computational complexities, making it well-suited for real-time communication applications where computational latency and power consumption must be adjusted in real time. Our FAJSCC is the first deepJSCC model that can dynamically adjust computational complexity with high performance.
@@
  • Computational Complexity vs. Performance: By varying the computational resources allocated to the encoder and decoder across different SNRs, we can identify which function is most sensitive to the variation of computational complexity and has the greatest impact on performance. Our experiment suggests that the decoder’s error correction capability is the most critical factor influencing performance, providing insight for future advancements aimed at maximizing the efficiency of deepJSCC. Note that FAJSCC is the first deepJSCC model that enables this analysis without modifying the architecture or scaling the model.
@@ II Problem Formulation
where $\mathcal{CN}(0,\sigma^{2})$ is the circularly symmetric complex Gaussian distribution with mean 00 and variance $\sigma^{2}$. In the case of the fast Rayleigh fading channel with noise power $\sigma^{2}$, the relationship between the $i$-th symbols of $\mathbf{y}$ and $\hat{\mathbf{y}}$ is given by
@@
${\textrm{SSIM}}(\mathbf{x},\hat{\mathbf{x}}):=l(\mathbf{x},\hat{\mathbf{x}})^{\alpha}c(\mathbf{x},\hat{\mathbf{x}})^{\beta}s(\mathbf{x},\hat{\mathbf{x}})^{\gamma},$
where hyperparameters $\alpha,\beta,\gamma$ weight the contributions of brightness similarity comparison $l(\mathbf{x},\hat{\mathbf{x}})$, contrast similarity comparison $c(\mathbf{x},\hat{\mathbf{x}})$, and structural similarity comparison $s(\mathbf{x},\hat{\mathbf{x}})$, respectively. Since SSIM considers multiple aspects that contribute to human perception, a higher SSIM value indicates better image quality as perceived by humans.
To measure the efficiency of the deepJSCC model, we evaluate computational resource usage in terms of floating point operations (FLOPs) and memory usage in bytes. FLOPs refer to the number of arithmetic operations involving floating-point numbers, such as additions and multiplications.

### v2

@@ Abstract
Recent advances in deep learning-based joint source–channel coding (deepJSCC) have substantially improved communication performance, but their high computational cost hinders practical deployment. Moreover, certain applications require the ability to dynamically adapt computational complexity. To address these issues, we propose a Feature Importance-Aware deepJSCC (FAJSCC) model for image transmission that is both computationally efficient and adjustable. FAJSCC employs axis-dimension specialized computation, which performs efficient operations individually for each spatial and channel axis, significantly reducing computational cost while representing features effectively. It further incorporates selective deformable self-attention, which applies self-attention only to selected and adaptively adjusted regions, leveraging the importance and relations of input features to efficiently capture complex feature correlations. Another key feature of FAJSCC is that the number of selected important areas can be controlled separately by the encoder and the decoder, depending on the available computational budget. It makes FAJSCC the first deepJSCC architecture to allow independent adjustment of encoder and decoder complexity within a single trained model. Experimental results show that FAJSCC achieves superior image transmission performance under various channel conditions while requiring less computational complexity than recent state-of-the-art models. Furthermore, experiments independently varying the encoder and decoder’s computational resources reveal, for the first time in the deepJSCC literature, that understanding the meaning of noisy features in the decoder demands the greatest computational cost. The code is publicly available at github.com/hansung-choi/FAJSCCv2.
@@ I Introduction
Despite its success in communication performance, existing deepJSCC research is increasingly suffering from growing computational complexity incurred by large neural networks, along with associated problems like high power consumption, hardware costs, and latency, which prohibit its practical deployment. To reduce such computational burden, one common approach is to apply model compression techniques. For instance, pruning [28, 29, 30] and low-rank decomposition [31] compress trained models, successfully reducing the number of parameters (i.e., computational load) at the expense of performance degradation and extra training efforts.
Beyond such compression-based strategies, another line of work focuses on designing lightweight models [32, 33, 34, 35, 36]. For example, reducing the intermediate deepJSCC feature size before passing through computationally intensive blocks [36] or performing computations only on a fixed subset of feature channels [32, 34] can lower the computational load. In addition, approaches that remove the least important features have been proposed [33, 37]. While these methods alleviate computational burden without requiring extra training, reducing or deleting features and restricting computations to limited channel regions degrade image transmission performance. Therefore, there is an urgent need for deepJSCC designs that can improve transmission performance while reducing computational cost compared to the current state-of-the-art models.
@@
Another important function that is often required in applications is the ability to dynamically control computational complexity within a single model. For instance, in surveillance applications, it is essential to transmit images of large surveillance areas at low computational complexity (i.e., low resolution) to ensure minimal delay and power consumption for long-term monitoring, while transmitting at high computational complexity (i.e., high resolution) when detailed analysis is needed. To address the varying computational needs at the application level, it is essential to dynamically manage the tradeoff between computational complexity and data communication accuracy. However, to the best of our knowledge, this issue is largely unstudied for image transmission in existing literature.
To address both efficiency and dynamic control over computational complexity, we propose feature importance-aware deepJSCC (FAJSCC). FAJSCC employs dimension-specialized computation strategies that enhance computational efficiency while achieving high transmission performance. The first strategy is axis-dimension specialized computation. A standard convolution over all channels, height, and width can be factorized into a depthwise convolution over the spatial dimensions (height and width) and a pointwise convolution over the channel dimension. This decomposition substantially reduces the number of operations: Instead of jointly processing all input and output channels with multi-channel kernels, the depthwise convolution handles each channel separately with each single-channel kernel, and the pointwise convolution mixes channels with 1×1 kernels. By exploiting this separation, FAJSCC achieves significantly lower computational cost while maintaining the feature processing capacity of standard convolution. To further enhance these operations, FAJSCC integrates spatial and channel attention mechanisms. The spatial attention emphasizes key features along the spatial axes, whereas the channel attention highlights important channels [38]. Accordingly, spatial attention is applied before depth-wise convolution and channel attention before point-wise convolution, maximizing feature processing for each axis. These lightweight attention operations exhibit linear computational complexity with respect to the input feature sizes. Our experiments demonstrate that even using only axis-dimension specialized computation can outperform the current state-of-the-art SwinJSCC [8] while using only half of the computational resources.
The second strategy is important-dimension specialized computation, which captures relations of complex features that axis-dimension operations may miss. To this end, FAJSCC employs a modified window self-attention mechanism [39]. The window self-attention mechanism shows highly effective feature processing ability by capturing relations of complex features for a given fixed window area in various semantic communication scenarios [40, 16, 8, 9, 41]. However, conventional window self-attention has two limitations: it cannot capture highly related features across different windows, and analyzing features of every window is computationally expensive. FAJSCC addresses these issues through a new attention mechanism, which we call selective deformable self-attention. The deformable self-attention adjusts self-attention regions based on the window and offsets computed from the input feature relation, enabling adaptive feature analysis beyond the fixed window.
Importantly, FAJSCC applies this deformable self-attention selectively based on feature importance. Some features are more informative (e.g., main objects), while others are less informative (e.g., background regions) [42]. Less informative features can be processed sufficiently by the preceding axis-dimension specialized operations alone. Moreover, simple images require only minimal computation to achieve high task performance [43, 44]. Inspired by these findings, FAJSCC applies this deformable self-attention only to the most important features rather than to all features. This selective enhancement strategy greatly reduces computational cost compared with applying self-attention across all windows. A portion of the saved computation is then reinvested to increase intermediate feature sizes, leading to richer representations. Although this reinvestment slightly raises the complexity, the increase is much smaller than the computational savings, so the overall cost remains significantly lower than applying self-attention to the entire feature set. As a result, FAJSCC achieves higher transmission performance by enhancing important feature processing while still maintaining reduced overall complexity. Unlike previous lightweight deepJSCC methods that reduce or remove features [36, 33, 37], which degrade performance, FAJSCC enlarges and selectively refines important features, which increase performance while reducing computational burden.
Finally, to further maximize efficiency, FAJSCC introduces a novel attention family tree that simultaneously extracts spatial and channel attention features, offsets, and feature importance while eliminating overlapping computations. Deformation offsets are derived from spatial-wise feature information, and spatial attention is generated from spatial-wise feature importance, which itself can be computed from the same spatial-wise feature information. Thus, redundant pre-processing steps can be removed by reusing spatial-wise feature information to get spatial-wise feature importance. Moreover, spatial-wise feature importance serves a dual purpose: identifying feature importance for selective self-attention and producing spatial attention features. Consequently, attentions, offsets, and feature importance are obtained efficiently without incurring significant computational overhead. Ablation studies confirm that the attention family tree achieves comparable or superior performance while reducing computational cost relative to extracting these features separately.
To adapt to varying computational budgets, we introduce the importance ratio $\gamma$, representing the fraction of features selected as important. When $\gamma=1$, all features are processed via deformable self-attention, whereas $\gamma=0$ means none are processed. For $0<\gamma<1$, the top $\gamma$ fraction of features, ranked by their measured importance, are selected and processed through deformable self-attention. By adjusting the importance ratio, FAJSCC enables controllable computational complexity within a single model. Previous resource-adjustable deepJSCC methods achieve complexity control by deleting features [33, 37], which can lead to significant performance degradation due to information loss. In contrast, FAJSCC maintains high performance because its selective feature enhancement strategy preserves all information.
Moreover, FAJSCC decouples computational complexity from the model architecture, allowing first independent evaluation of the impact of encoder and decoder complexity on performance. It is important to note that typical previous deepJSCC models have fixed and symmetric computational complexity at encoder and decoder [6, 8, 13] or the decoder’s computational complexity is inherently constrained by feature deletion at the encoder [33, 37]. As a result, such a discussion has not been possible in previous models. Recall that the encoder compresses images and adds redundancy to protect data against channel noise, while the decoder interprets the received noisy signal and (re)generates the transmitted images. A natural question arises: Which component requires the most computational complexity, and under what conditions? By independently varying the importance ratios for the encoder and decoder across different signal-to-noise ratios (SNRs), we draw two key conclusions: 1) the decoder requires higher computational complexity than the encoder, and 2) the decoder’s noisy-signal perception demands even greater complexity when the SNR is low. To the best of our knowledge, this is the first detailed analysis of computational complexity in deepJSCC.
As a consequence of our design, our FAJSCC achieves higher image transmission performance than other deepJSCC models, even with lower computational complexity compared to the previous state-of-the-art (SOTA) model [8]. Our FAJSCC also shows higher efficiency than the recent light-weight model [32]. Our main contributions can be summarized as follows.
  • Efficient Computation: We propose axis-dimension specialized computation and selective deformable self-attention to significantly reduce computational cost while maintaining high transmission performance. By processing only the most important features with deformable self-attention and efficiently reusing attention-related information through the attention family tree, FAJSCC achieves superior image reconstruction compared to existing SOTA models with lower computational requirements.
  • Adjustable Computation: We introduce the importance ratio to control the fraction of features processed by deformable self-attention, enabling FAJSCC to dynamically adjust computational complexity within a single model. Unlike previous methods that delete features to fit computational resources, FAJSCC preserves all information and allows independent adjustment of encoder and decoder complexity, making it possible to allocate computation according to available resources without compromising performance.
@@
  • Computational Complexity vs. Performance: By varying the computational resources allocated to the encoder and decoder across different SNRs, we can identify which function is most sensitive to the variation of computational complexity and has the greatest impact on performance. Our experiment suggests that the decoder’s noisy-signal perception is the most critical factor influencing performance, providing insight for future advancements aimed at maximizing the efficiency of deepJSCC. Note that FAJSCC is the first deepJSCC model that enables this analysis without modifying the architecture or scaling the model.
@@ II Problem Formulation
where $\mathcal{CN}(0,\sigma^{2})$ is the circularly symmetric complex Gaussian distribution with mean 0 and variance $\sigma^{2}$. In the case of the fast Rayleigh fading channel with noise power $\sigma^{2}$, the relationship between the $i$-th symbols of $\mathbf{y}$ and $\hat{\mathbf{y}}$ is given by
@@
${\textrm{SSIM}}(\mathbf{x},\hat{\mathbf{x}}):=l(\mathbf{x},\hat{\mathbf{x}})^{p}c(\mathbf{x},\hat{\mathbf{x}})^{q}s(\mathbf{x},\hat{\mathbf{x}})^{r},$
[…]
To measure the efficiency of the deepJSCC model, we evaluate computational resource usage in terms of floating-point operations (FLOPs) and memory usage in bytes. FLOPs refer to the number of arithmetic operations involving floating-point numbers, such as additions and multiplications.